# Grid-Based Continuous Normal Representation for Anomaly Detection

### v1

@@ Abstract
There have been significant advancements in anomaly detection in an unsupervised manner, where only normal images are available for training. Several recent methods aim to detect anomalies based on a memory, comparing the input and the directly stored normal features (or trained features with normal images). However, such memory-based approaches operate on a discrete feature space implemented by the nearest neighbor or attention mechanism, suffering from poor generalization or an identity shortcut issue outputting the same as input, respectively. Furthermore, the majority of existing methods are designed to detect single-class anomalies, resulting in unsatisfactory performance when presented with multiple classes of objects. To tackle all of the above challenges, we propose GRAD, a novel anomaly detection method for representing normal features within a “continuous” feature space, enabled by transforming spatial features into coordinates and mapping them to continuous grids. Furthermore, we carefully design the grids tailored for anomaly detection, representing both local and global normal features and fusing them effectively. Our extensive experiments demonstrate that GRAD successfully generalizes the normal features and mitigates the identity shortcut, furthermore, GRAD effectively handles diverse classes in a single model thanks to the high-granularity global representation. In an evaluation using the MVTec AD dataset, GRAD significantly outperforms the previous state-of-the-art method by reducing 65.0% of the error for multi-class unified anomaly detection. The project page is available at https://tae-mo.github.io/grad/.
@@
Anomaly detection Continuous normal space Grid representation
Figure 1: Conceptual diagram and qualitative results of existing methods and ours. (a) and (b) use single and multiple normal features in a discrete space, respectively, while our method (c) exploits continuous feature space. We visualize the anomaly detection process with the normal (navy) and abnormal (red) patches of the top-left reference image. ‘Pred.’ indicates the prediction based on the disparity, and wrong predictions are marked as (X) with red color. We present the reconstruction results based on the reference abnormal images.
@@ 1 Introduction
With the recent advances in deep neural networks, anomaly detection (AD) has been applied for a wide range of applications such as manufacturing industry [20, 30], video surveillance [33, 35], and medical imaging [34, 38]. Despite its success, there are still several limitations hindering the broader applicability across many practical scenarios. In particular, one major bottleneck is the collection of a sufficient amount of anomalous data, which is scarce by definition of an anomaly. Furthermore, many AD systems often require considerable effort for pixel-wise labeling of ‘normal’ versus ‘abnormal’ data. Due to these challenges, there is a growing interest in developing methods in an unsupervised manner, which have demonstrated very promising results even when trained solely with normal data  [30].
@@
As notable examples, PatchCore [30], PaDim [8], and SPADE[7] proposed using additional memory that directly stores normal features (or distributions). During inference, these memory-based methods detect anomalies based on the distance between the testing input and its nearest neighbor in their memory, as shown in Fig. 1(a). While showing promising performance, these methods necessitate storing a wide array of diverse normal features in memory, resulting in high space complexity and resource-intensive search operations. Moreover, these methods are often ineffective in identifying the characteristics of global anomalies due to their diversity, resulting in suboptimal AD performance (see Tab. 1).
@@
Another line of work [13, 25, 15] focuses on producing generalized normal features. Unlike the aforementioned approaches that use the nearest neighbor technique, these methods combine multiple normal features from the memory using an attention mechanism (i.e., referring to multiple discrete features), given a normal or abnormal input (Fig. 1(b)). They assume that the model always generates normal features, regardless of whether the inputs are normal or abnormal, thus anomalous regions can be detected based on the disparity between the inputs and outputs. Because these models gather diverse normal features from the memory via attention, they have exhibited increased robustness to test data, leading to improved generalization performance. However, such strength may turn into a drawback when testing abnormal inputs. If these inputs can be reconstructed using a combination of normal features, the model might ultimately generate an output that is identical to the abnormal input. This issue, termed as an identity shortcut (IS) by UniAD [40], prevents the models from detecting anomalies due to the lack of disparity between the abnormal input and produced output.
@@
Furthermore, a significant limitation exists in most of the approaches discussed earlier, as they are primarily designed to handle only one class of objects per model. When these methods are extended to multi-class scenarios, where a single model handles multiple classes, a significant performance drop is observed [40], even with state-of-the-art methods. To mitigate this limitation, memory-based methods may incorporate sufficient memory to accommodate multi-class normality, yet this simultaneously increases memory consumption search latency. Moreover, attention-based methods experience more severe challenges with the IS problem, as the greater number and diversity of aggregated features tend to more easily represent anomalies. These challenges necessitate training distinct models for each class, which increases training complexity, memory usage, computational overheads, and even data preparation efforts in a practical implementation.
To address all of the aforementioned issues, we propose Grid Representation for Anomaly Detection (GRAD), which adopts a “continuous” normal feature space instead of a discrete one (Fig. 1(c)). The capability of a continuous feature space to represent infinitely fine granularity allows for a certain amount of variation in normal features, thereby mitigating the weak generalization issue. Also, as our approach does not rely on combining or aggregating multiple features, it prevents the creation of entirely new features (unseen anomalies in our context), thus helping to avoid the IS problem.
We create the continuous space using a grid structure, successfully adopted in the neural fields domain for signal representation, which functions by taking coordinates as input and outputting corresponding quantities. However, given that the input to the grid for reconstructing a normal feature is a spatial feature from an input image, rather than coordinates as in conventional grids, we require a specially designed framework optimized for handling feature-based inputs. In light of this requirement, we propose a novel grid-based AD framework, where input spatial features are transformed into low-dimensional coordinates and then mapped onto the grid structure. To enable grids to represent normal spatial features in a continuous space, we utilize both local and global grids, each designed to represent different aspects of normal features. Then, we seamlessly fuse these two grids to effectively capture coarse-to-fine anomalies. In addition, we introduce coordinate jittering, which allows the grids to learn more generalized representations by updating more grid values with each coordinate. Furthermore, we minimize the false detections (i.e., false positives) through the feature refinement process, which forces the normal regions of fused output to be the same as the original input.
@@
GRAD not only enhances AD performance by effectively addressing weak generalization and IS issues but also provides several additional advantages. As its global representation can capture the structural characteristics of normal images, GRAD is provided with the capability for identifying anomalies in more diverse classes, each characterized by distinct structures. In terms of computational and parameter efficiency, GRAD avoids the tremendous computational overhead associated with searching every feature as in the existing methods (Fig. 1(a),(b)), since GRAD acquires a single feature directly from the continuous space. Furthermore, the high granularity of continuous space enables compact representation, resulting in high parameter efficiency. To the best of our knowledge, this is the first work to apply the continuous feature space to effectively represent normal features for AD.
@@
In the extensive experiments, we demonstrate the superiority of continuous normal space over the existing discrete spaces in terms of accuracy and efficiency (both for computation and parameter). Furthermore, experimental results on the MVTec AD dataset [3] show that GRAD achieves state-of-the-art performance in a unified setting (multi-class AD with a single model) for both detection and localization, which even outperforms the state-of-the-art models trained for each respective class. With the comprehensive analysis, we demonstrate GRAD is an effective solution for AD, overcoming the limitations of existing methods.
@@ 2.1 Unsupervised Anomaly Detection
Other methods leverage auxiliary memory to retain normal features, where they are classified into the following two categories: reference to a single discrete feature (Fig. 1(a)) and reference to a combination of discrete features (Fig. 1(b)). The former methods [7, 28, 8, 30] detect anomalies by measuring the distance between the input and stored features (or feature distributions) extracted by a pre-trained network. For instance, PatchCore [30] and SPADE [7] are designed to store the representative normal features in a memory bank and use the nearest neighbor search for anomaly scoring. However, the above methods present a significant challenge to represent features that are not already stored in their memory. Therefore, when faced with a complex and diverse range of inputs, they result in limited performance or require tremendous memory usage to achieve satisfactory performance.
Moreover, an increase in the number of stored features can significantly delay the time to reference all the stored ones, further hampering their effectiveness. The latter methods [13, 15, 25] employ attention-like techniques to take a weighted sum of all normal features in the discrete space based on their similarity to the input. While these approaches exhibit superior generalization capabilities compared to the former methods, they generalize not only to normal features but also to abnormal features, using combinations of all features in the memory. This causes the input anomalies to be reconstructed, coined as the IS problem, reducing the AD performance by hindering the model from recognizing the disparity, as depicted in Fig. 1(b).
@@ 2.3 Grid Representation
Figure 2: (a) The detailed architecture of GRAD and (b) visualization of coordinate jittering. The input $x$ is firstly transformed into pixel-wise and feature-wise coordinates. After the normal features are sampled from local and global representations, they are fused by CNN blocks. The final reconstruction is acquired through the proposed feature refinement process.
3 GRAD
Background. To facilitate the readers to understand GRAD, we first describe the grid operation. A grid is trained as a function of coordinates with infinite resolution, outputting coordinate-corresponding features. The output feature in infinite resolution is aggregated by nearby features in the grid, based on the distance between the coordinate of the input and neighboring features. For example, when we take 1D grid sampling $\phi(\cdot;G):\mathbb{R}\rightarrow\mathbb{R}^{C}$, the output feature with channel $C$ is interpolated by neighboring values of the 1D grid $G\in\mathbb{R}^{R\times C}$, which is mathematically formulated as follows:
@@ 3 GRAD
Overview. The motivation of our work is to effectively represent the normal features in continuous space using the grid operation, distinct from the discrete spaces. In an unsupervised manner, GRAD detects anomalous images and regions based on the discrepancy between the input feature and output normal feature, as described in Fig. 2(a). Therefore, the primary objective of GRAD is to effectively retain the normal components (e.g., shapes or textures) of the original input feature while eliminating any anomalies presented within the feature.
To this end, we represent normal features in the continuous space during the training phase, coined as normal representation, which is used for replacing abnormal features in the testing phase. We describe how GRAD represents normal features in the continuous space and acquires the output feature $\hat{x}$, based on the input feature $x$ extracted from a pre-trained backbone, where $x$, $\hat{x}\in\mathbb{R}^{C\times H\times W}$, and $C,H,W$ is the channel, height, and width of the feature, respectively.
@@ 3.1 Normal Representation
The fundamental concept of GRAD is to transform the input feature into specific coordinates of continuous values, which are subsequently mapped to feature grids. In particular, we design to represent the normal features from local and global perspectives. By combining the distinctive features from each perspective, the resulting feature can provide a strong representation of the input, capturing both fine-grained details as well as broader overall structures.
@@
In addition, for an abnormal patch, GRAD finds a normal feature that is the most representative of the abnormal patch based on the reduced coordinates. As the grid has never been exposed to abnormal features during training, it is unable to represent abnormal features by interpolating nearby normal features. This is the core idea of how we can effectively resolve the identity shortcut (IS) issue frequently found in the existing methods that aggregate numerous features based on similarities using attention mechanisms [13, 25].
@@
[…]
where $\mathsf{concat}(\cdot,\cdot)$ denotes the concatenation of two features along with the channel axis. By fusing the local and global representation, GRAD can represent normal features from fine-grained details to broader contexts, resulting in higher performance compared to the cases using only either of them (see ablation study in Sec. 4.4).
@@ 3.3 Training and Inference
Figure 3: Visualization of GRAD’s pipeline. Each marker in (a) corresponds to the patch on the left image that has the same number and color. Each marker in (b) corresponds to a single image from the test dataset, where different colors represent distinct classes, and circles and triangles denote the normal and abnormal images, respectively. ‘Dim 1’ and ‘Dim 2’ are the two dimensions of 2D grids.
Persp. Method #Entry Detection Localization
Local VQ 64 97.0±0.85 96.0±0.06
256 97.4±0.19 96.1±0.24
Attention 64 95.9±0.36 96.4±0.06
256 95.5±0.26 96.2±0.06
GRAD 64 98.0±0.05 97.1±0.07
Global VQ 16 78.9±1.28 86.7±1.93
64 76.9±0.73 86.1±0.49
Attention 16 79.8±1.22 88.0±0.27
64 80.1±2.56 87.8±1.80
GRAD 16 93.0±0.34 95.7±0.04
Table 1: Performance evaluation of the different feature space in a unified setting. #Entry denotes the number of features in each feature space and Persp. indicates the perspective (local or global).
@@ 4.1 Experimental Setup
We used MVTec AD [3] and VisA [44] datasets, which are representative datasets for real-world unsupervised AD. We evaluated the performance of anomaly detection by the Area Under the Receiver Operator Curve (AUROC). Following previous studies, we computed the class-average AUROC for detection and pixel-wise AUROC for localization. We implemented GRAD in the PyTorch framework, and we used the NVIDIA A5000 GPU for all evaluations. We trained our models for 100 epochs, thrice with different seeds (0,1,2), with a batch size of 64. We describe the detailed implementation of GRAD in Appendix 0.A.
@@
4.2 Effectiveness of Continuous Feature Space
Improved performance. To assess the efficacy of the continuous normal space, we implemented two baselines with discrete feature spaces under the same overall detection framework of GRAD as follows: 1) referring to a single feature from discrete space (Fig. 1(a)) through vector quantization (VQ), and 2) referring to a combination of multiple discrete features (Fig. 1(b)) with an attention module. As shown in Tab. 1, GRAD, providing a continuous normal feature space, outperforms the other baselines for both local and global representation. When we expand the feature space for local representation, the attention shows the performance drop, suffering from a more severe IS. Although VQ shows performance improvement with a larger feature space, it still falls short of GRAD even with the quadrupled feature space. Furthermore, the baselines consistently underperform in global representation, indicating their inability to represent structural information of the entire feature.
Visualization of coordinates. Although the quantitative results above clearly demonstrate the effectiveness of the continuous normal representation of GRAD, we additionally visualize the generated and mapped coordinates in Fig. 3. The normal and abnormal areas with a similar local characteristic are mapped at a near distance in the local feature space (e.g., the patch $317$ and $485$ in Fig. 3(a)). Similarly, the global coordinates of the two input images are mapped to almost the same location at the global feature space (Fig. 3(b)). These results indicate that the model successfully learns to generate coordinates corresponding to each input feature, and the local and global grids can represent the normal features from each perspective effectively.
Class US [4] PaDiM [8] MKD [32] DRAEM [41] RD4AD [9] PatchCore [30] UniAD [40] HVQ-T [23] GRAD (Ours)

### v3

# Continuous Memory Representation for Anomaly Detection

@@ Abstract
There have been significant advancements in anomaly detection in an unsupervised manner, where only normal images are available for training. Several recent methods aim to detect anomalies based on a memory, comparing or reconstructing the input with directly stored normal features (or trained features with normal images). However, such memory-based approaches operate on a discrete feature space implemented by the nearest neighbor or attention mechanism, suffering from poor generalization or an identity shortcut issue outputting the same as input, respectively. Furthermore, the majority of existing methods are designed to detect single-class anomalies, resulting in unsatisfactory performance when presented with multiple classes of objects. To tackle all of the above challenges, we propose CRAD, a novel anomaly detection method for representing normal features within a “continuous” memory, enabled by transforming spatial features into coordinates and mapping them to continuous grids. Furthermore, we carefully design the grids tailored for anomaly detection, representing both local and global normal features and fusing them effectively. Our extensive experiments demonstrate that CRAD successfully generalizes the normal features and mitigates the identity shortcut, furthermore, CRAD effectively handles diverse classes in a single model thanks to the high-granularity continuous representation. In an evaluation using the MVTec AD dataset, CRAD significantly outperforms the previous state-of-the-art method by reducing 65.0% of the error for multi-class unified anomaly detection. Our project page is available at https://tae-mo.github.io/crad/.
@@
Anomaly detection Continuous memory representation
footnotetext: * Equal contribution.
Corresponding authors.
Figure 1: Conceptual diagram and qualitative results of existing methods and ours. (a) and (b) use single and multiple normal features in a discrete memory, respectively, while our method (c) exploits continuous feature memory. We visualize the anomaly detection process with the normal (navy) and abnormal (red) patches of the top-left reference image. ‘Pred.’ indicates the prediction based on the disparity, and wrong predictions are marked as (X) with red color. We present the reconstruction results based on the reference abnormal images.
@@ 1 Introduction
With the recent advances in deep neural networks, anomaly detection (AD) has been applied for a wide range of applications such as manufacturing industry [19, 28], video surveillance [31, 33], and medical imaging [32, 35]. Despite its success, there are still several limitations hindering the broader applicability across many practical scenarios. In particular, one major bottleneck is the collection of a sufficient amount of anomalous data, which is scarce by definition of an anomaly. Furthermore, many AD systems often require considerable effort for pixel-wise labeling of ‘normal’ versus ‘abnormal’ data. Due to these challenges, there has been a growing interest in developing methods in an unsupervised manner where we train AD models solely with normal data [28].
@@
As notable examples, PatchCore [28], PaDiM [8], and SPADE[7] proposed using additional memory that directly stores normal features (or distributions). During inference, these memory-based methods detect anomalies based on the distance between the testing input and its nearest neighbor in their memory, as shown in Fig. 1(a). While showing promising performance, these methods require storing a wide array of diverse normal features in memory, resulting in high space complexity and resource-intensive search operations. Moreover, these methods are often ineffective in identifying the characteristics of global anomalies due to their diversity, leading to suboptimal AD performance (see Sec. 4.2).
@@
Another line of work [13, 23, 14] focuses on producing generalized normal features. Unlike the aforementioned approaches that use the nearest neighbor technique, these methods combine multiple normal features from the memory using an attention mechanism (i.e., referring to multiple discrete features), given a normal or abnormal input (Fig. 1(b)). They assume that the model always generates normal features, regardless of whether the inputs are normal or abnormal, thus anomalous regions can be detected based on the disparity between the inputs and outputs. Because these models gather diverse normal features from the memory via attention, they have exhibited increased robustness to test data, leading to improved generalization performance. However, such strength may turn into a drawback when testing abnormal inputs. If these abnormal inputs can be reconstructed using a combination of normal features, the model could potentially generate outputs that are identical to the abnormal inputs. This issue, referred to as an identity shortcut (IS) by UniAD [37], prevents the models from detecting anomalies due to the minimal difference between the abnormal input and produced output.
@@
Furthermore, a significant limitation exists in most of the approaches discussed earlier, as they are primarily designed to handle only one class of objects per model. When these methods are extended to multi-class scenarios, where a single model handles multiple classes, a significant performance drop has been observed [37], even with state-of-the-art methods. To mitigate this limitation, memory-based methods may incorporate sufficient memory to accommodate multi-class normality, yet this simultaneously increases memory consumption and search latency. Moreover, attention-based methods experience more severe challenges with the IS problem, as the greater number and diversity of aggregated features tend to more easily represent anomalies. These challenges necessitate training distinct models for each class, which increases training complexity, memory usage, computational overheads, and even data preparation efforts in a practical implementation.
To address all of the issues above, we propose a novel continuous memory representation for anomaly detection (CRAD), where we use an external grid-based representation for normal features (Fig. 1(c)). Given that the input to the grid is a spatial feature from an image rather than coordinates required for conventional grids, we need a specially designed framework for handling feature-based inputs. In light of this requirement, we transform input spatial features into low-dimensional coordinates, based on which we interpolate neighboring normal features in the grid. This continuous memory, unlike other discrete counterparts, allows for instant ($\mathcal{O}(1)$ time complexity) retrieval of the normal feature, while the interpolation technique mitigates the weak generalization issue. Also, as our approach does not rely on innumerable features across the entire memory, it reduces the risk of generating entirely new features (unseen anomalies in our context), thus helping to avoid the IS problem. These advantages are even stronger in the multi-class scenario, where a larger memory space would be needed, by avoiding the tremendous computation associated with searching or aggregating every feature.
Deploying the continuous memory, we additionally include specific designs on CRAD tailored for AD. CRAD incorporates two distinct continuous memories to represent normal features from both a local and global perspective. Through the integration of these representations, CRAD adeptly identifies coarse-to-fine anomalies. Furthermore, we implement coordinate jittering to enhance the generalization capability of the grid, facilitating the update of a broader range of grid values with each input coordinate. A feature refinement process further improves CRAD, minimizing false detections (i.e., false positives) by ensuring that the normal regions in the fused output remain consistent with the original input.
@@
CRAD not only addresses weak generalization and IS issues but also provides several additional advantages. Whereas discrete memories struggle to represent global features (e.g., an entire input feature), our continuous memory successfully captures their structural characteristics (Sec. 4.2). Therefore, it enables identifying anomalies across a wider range of classes, each featured by distinct structures. Furthermore, different from discrete memories with limited entries, CRAD represents an infinite number of normal features in the continuous memory. Thus, CRAD achieves high performance with compact memory, resulting in high parameter efficiency. To the best of our knowledge, this is the first work leveraging continuous memory to effectively represent normal features for AD.
@@
In the extensive experiments, we demonstrate the superiority of continuous memory over the existing discrete spaces in terms of accuracy and efficiency (both for computation and parameter). Furthermore, experimental results on the MVTec AD dataset [3] show that CRAD achieves state-of-the-art performance in a unified setting (multi-class AD with a single model) for anomaly detection, which even outperforms the state-of-the-art models trained for each respective class. With the comprehensive analysis, we demonstrate CRAD is an effective solution for AD, overcoming the limitations of existing methods.
@@ 2.1 Unsupervised Anomaly Detection
Other methods leverage auxiliary memory to retain normal features, where they are classified into the following two categories: reference to a single discrete feature (Fig. 1(a)) and reference to a combination of discrete features (Fig. 1(b)). The former methods [7, 26, 8, 28] detect anomalies by measuring the distance between the input and stored features (or feature distributions) extracted by a pre-trained network. For instance, PatchCore [28] and SPADE [7] are designed to store the representative normal features in a memory bank and use the nearest neighbor search for anomaly scoring. However, the above methods present a significant challenge to represent features that are not already stored in their memory. Therefore, when faced with a complex and diverse range of inputs, they result in limited performance or require tremendous memory usage to achieve satisfactory performance. Moreover, an increase in the number of stored features can significantly delay the time to reference all the stored ones, further hampering their effectiveness.
The latter methods [13, 14, 23] employ attention-like techniques to take a weighted sum of all normal features in the discrete space based on their similarity to the input. While these approaches exhibit superior generalization capabilities compared to the former methods, they generalize not only to normal features but also to abnormal features, using combinations of all features in the memory. This causes the input anomalies to be reconstructed, coined as the IS problem, reducing the AD performance by hindering the model from recognizing the disparity, as depicted in Fig. 1(b).
@@ 2.3 Grid Representation
Figure 2: (a) The detailed architecture of CRAD and (b) visualization of coordinate jittering. The input $x$ is firstly transformed into pixel-wise and feature-wise coordinates. After the normal features are sampled from local and global representations, they are fused by CNN blocks. The final reconstruction is acquired through the proposed feature refinement process.
3 CRAD
Background. To help the readers understand CRAD, we first describe the grid operation. A grid is trained as a function of coordinates with infinite resolution, outputting coordinate-corresponding features. The output feature in infinite resolution is aggregated by nearby features in the grid, based on the distance between the coordinate of the input and neighboring features. For example, when we take 1D grid sampling $\phi(\cdot;G):\mathbb{R}\rightarrow\mathbb{R}^{C}$, the output feature with channel $C$ is interpolated by neighboring values of the 1D grid $G\in\mathbb{R}^{R\times C}$, which is mathematically formulated as follows:
@@ 3 CRAD
Overview. The motivation of our work is to effectively represent the normal features in continuous memory using the grid operation, distinct from the discrete memories. In an unsupervised manner, CRAD detects anomalous images and regions based on the discrepancy between the input feature and output normal feature, as described in Fig. 2(a). Therefore, the primary objective of CRAD is to effectively retain the normal components (e.g., shapes or textures) of the original input feature while eliminating any anomalies presented within the feature.
To this end, we represent normal features in the continuous memory during the training phase, coined as normal representation, which is used for replacing abnormal features in the testing phase. We describe how CRAD represents normal features in the continuous memory and acquires the output feature $\hat{x}$, based on the input feature $x$ extracted from a pre-trained backbone, where $x$, $\hat{x}\in\mathbb{R}^{C\times H\times W}$, and $C,H,W$ are the channel, height, and width of the feature, respectively.
@@ 3.1 Normal Representation
The fundamental concept of CRAD is to transform the input feature into specific coordinates of continuous values, which are subsequently mapped to feature grids. In particular, we design to represent the normal features from local and global perspectives. By combining the distinctive features from each perspective, the resulting feature can provide a strong representation of the input, capturing both fine-grained details as well as broader overall structures.
@@
In addition, for an abnormal patch, CRAD finds a normal feature that is the most representative of the abnormal patch based on the reduced coordinates. As the grid has never been exposed to abnormal features during training, it is unable to represent abnormal features by interpolating nearby normal features. This is the core idea of how we can effectively resolve the identity shortcut (IS) issue frequently found in the existing methods that aggregate numerous features based on similarities using attention mechanisms [13, 23].
@@
where $\mathsf{concat}(\cdot,\cdot)$ denotes the concatenation of two features along with the channel axis. By fusing the local and global representation, CRAD can represent normal features from fine-grained details to broader contexts, resulting in higher performance compared to the cases using only either of them (see ablation study in Sec. 4.4).
@@ 4.1 Experimental Setup
We used MVTec AD [3] and VisA [41] datasets, which are representative datasets for real-world unsupervised AD. We evaluated the performance of anomaly detection by the Area Under the Receiver Operator Curve (AUROC). Following previous studies, we computed the class-average AUROC for detection and pixel-wise AUROC for localization. We implemented CRAD in the PyTorch framework, and we used the NVIDIA A5000 GPU for all evaluations. We trained our models for 50 epochs, thrice with different seeds (0,1,2), with a batch size of 64. We describe the detailed implementation of CRAD in the supplementary materials.
@@
4.2 Effectiveness of Continuous Memory Representation
Improved performance. To assess the efficacy of the continuous memory, we implemented two baselines with discrete memories under the same overall detection framework of CRAD as follows: 1) referring to a single feature from discrete space (Fig. 1(a)) through vector quantization (VQ), and 2) referring to a combination of multiple discrete features (Fig. 1(b)) with an attention module. As shown in Sec. 4.2, CRAD, providing a continuous memory, outperforms the other baselines for both local and global representation. When we expand the memory size for local representation, the attention shows the performance drop, suffering from a more severe IS.
Table 1: Performance evaluation of the different feature memories in the unified setting. #Entry denotes the number of features in each memory and Persp. indicates the perspective (local or global).
Persp. Method #Entry Detection Localization
Local VQ 64 96.9±0.65 96.1±0.05
256 97.8±0.23 96.0±0.08
Attention 64 95.9±1.1 96.2±0.25
256 93.9±1.8 95.1±0.76
CRAD 64 98.6±0.07 97.5±0.04
Global VQ 16 81.0±0.97 89.9±0.42
64 82.2±0.56 91.4±1.0
Attention 16 77.9±2.3 86.7±3.4
64 82.1±0.54 90.6±0.19
CRAD 16 92.3±0.60 95.7±0.17
Although VQ shows performance improvement with larger memory entries, it still falls short of CRAD even with the quadrupled feature space. Furthermore, the baselines consistently underperform in global representation, indicating their inability to represent structural information of the entire feature.
Visualization of coordinates. Although the quantitative results above clearly demonstrate the effectiveness of the continuous normal representation of CRAD, we additionally visualize the generated and mapped coordinates in Fig. 3. The normal and abnormal areas with a similar local characteristic are mapped at a near distance in the local feature space (e.g., the patch $317$ and $485$ in Fig. 3(a)). Similarly, the global coordinates of the two input images are mapped to almost the same location at the global feature space (Fig. 3(b)). These results indicate that the model successfully learns to generate coordinates corresponding to each input feature, and the local and global grids can represent the normal features from each perspective effectively.
Figure 3: Visualization of CRAD’s pipeline. Each marker in (a) corresponds to the patch on the left image that has the same number and color. Each marker in (b) corresponds to a single image from the test dataset, where different colors represent distinct classes, and circles and triangles denote the normal and abnormal images, respectively. ‘Dim 1’ and ‘Dim 2’ are the two dimensions of 2D grids.
Class US [4] PaDiM [8] MKD [30] DRAEM [38] RD4AD [9] PatchCore [28] UniAD [37] HVQ-T [21] CRAD (Ours)